OpenStudy (calculusxy):

Need someone to check my answers for a conditional statement, converse, and contrapositive statement! MEDAL!!!

OpenStudy (calculusxy):

Statement: If x = 1, then \(x^2\) = x. a. Is it true? `Yes because if you have (1)^2 = 1 x 1 = 1`

OpenStudy (calculusxy):

b. Now state the converse. Is this true? Explain. Draw a picture if necessary. If x^2 = x, then x = 1. False because x can not only equal to 1, but it can also equal to 0. Work: 0^2 = 0 x 0 = 0

OpenStudy (calculusxy):

c. Write the contra-positive statement. If x doesn't equal to 1, then x^2 doesn't equal to x.

OpenStudy (calculusxy):

Are all of my answers correct?

pooja195 (pooja195):

I agree with your answers.

OpenStudy (calculusxy):

Thank you!